Dimi has the following utility for goods x and y:

u(x,y) = x^2 + 2x + 1 - y

The marginal rate of substitution between the two goods is...


a) 2

b) 2x-2

c) 2x+2

d) -2x+2


1
Expert's answer
2021-05-24T08:58:56-0400

Solution:

Marginal rate of substitution (MRS) = "\frac{Change \;in\; Good \;Y}{Change \;in\; Good \;X} = \frac{MU_{X} }{MU_{Y}}"

MRSx,y = "\frac{MU_{X} }{MU_{Y}}"  

First derive MUx:


MUx  = "\frac{\partial U} {\partial X} = 2x +2"


MUy  = "\frac{\partial U} {\partial Y} = -1"

Substitute to derive MRSx,y:

MRSx,y =  "\frac{MU_{X} }{MU_{Y}} = \frac{2x +2 }{-1} = -2x - 2"


MRSx,y ="-2x - 2"